# SETA: Semantic-Aware Token Augmentation for Domain Generalization

### v1

@@ III-C Enhancing Shape Bias by Texture Noise
While edge representations encompass both global shape edges and local texture edges, directing the model to learn edge representations may still lead to sensitivity to local texture edges. To tackle this issue, we propose the Shape-disrupted Token Shuffling (STS) module to generate local edge noise, which is incorporated into object edge maps for potentially enhancing model learning of shape features.
Shape-disrupted token shuffling. Motivated by the insight that holistic shapes could be discerned through the relative positions of edge tokens [17, 46, 23], we introduce texture noise using shape-disrupted samples by shuffling tokens within the samples. For the input representation $z_{j}\in\mathbb{R}^{N\times C}$, we employ ${\rm Shuffle}(\cdot)$ function in token dimension:
@@
where $\lambda$ is a random weight sampled from the Beta distribution ${\rm Beta}(\alpha,\beta)$ with $\alpha,\beta\in(0,\infty)$ being two hyper-parameters. The Mixup version of SETA can introduce bote style noise and local edge noise into the edge-related tokens, thereby encouraging the model to learn global semantic features from the edge representations. 2) For the CutMix version, we generate the token-level binary mask $\mathcal{M}^{\lambda}\in\mathbb{R}^{N\times C}$ with the element being set to $1$ if the corresponding elemnet in $E(\hat{z}_{i})$ is in the top $\lambda\sim{\rm Beta}(\alpha,\beta)$ percentage elements. The augmented representation $\widetilde{z}_{i}$ is:
$\widetilde{z}_{i}=\mathcal{M}^{\lambda}\odot z_{i}+(1-\mathcal{M}^{\lambda})\odot z_{j}^{s}.$ (8)
@@
In practice, for each mini-batch, we randomly select half of the samples to conduct the SETA augmentation. Within this process, the Mixup variant of SETA is applied with a probability of $p_{\text{mode}}$, while in other cases, the CutMix version of SETA is employed. In this way, the SETA could generate diverse augmentation for the edge-related tokens while preserving global semantic information, thus effectively enhancing the shape bias of the model. During inference, SETA is closed as conventional augmentation-based DG methods [14, 19].
@@
Two stylized variants of SETA. To adequately enhance model robustness to domain shifts, we further explore two stylized versions of SETA. As our SETA primarily relies on texture noise extracted from shape-disrupted neural samples, the diversity of styles within the augmented samples is still limited. To address this limitation, we have incorporated the two SOTA DG style augmentation method and devised two SETA variants, specifically SETA-S-DSU (combined with DSU [14]), which introduces statistical diversity, and SETA-S-ALOFT (combined with ALOFT [13]), which introduces frequency diversity. In this context, when generating token-shuffled samples with disrupted shapes, we also apply style augmentations (i.e., DSU or ALOFT) to these samples. The stylized variants could concurrently amplify the diversity of both style and texture edges, thus providing a more comprehensive simulation of potential domain shifts.
@@
In summary, we propose a token-level feature augmentation designed to encourage the shape bias of the model by introducing local edge noise. Note that our SETA diverges significantly from prior augmentation-based DG works [13, 14, 19, 33], which solely perturb style information while keeping other features unchanged. Since these existing methods ignore spurious local edges within the remaining features, the model inevitably learns domain-specific features and fails to capture shape information effectively. In contrast, our method proposes to interpolate edge-pass shape-preserved representations from the original image with shape-disrupted local edge representations from another image. This method compels the model not only to identify edges but also to distinguish between relevant and spurious ones. As a result, our method enhances the model learning of global shape information and promotes its ability to extract domain-invariant features. Besides, the computational overhead of SETA is minimal, involving only a few matrix operations. The components within SETA introduce no additional parameters or inference time, resulting in negligible training overhead, e.g., increasing only $0.6\%$ ($0.154s$ vs. $0.153s$) in time for each training step from the baseline on GFNet-H-Ti.
@@ III-D Theoretical Insight of SETA
Proof. Given the source dataset $D_{s}$, feature augmentation creates sensitive data to augment the training data, denoted as $D_{s}^{aug}$. Recall that Lemma $1$ provides a generalization error bound between two different distributions. For $h_{aug}$ trained on the augmented source dataset $D_{s}^{aug}$, the following inequality holds with a probability of at least $1-\delta$:
@@
Theorem $1$ indicates that the upper bound of generalization risk on target domain depends primarily on two terms: 1) $\varepsilon(D_{s},D_{s}^{aug})$ that denotes the consistency between the risk of the model on clean and augmented data; 2) $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ that presents the distance between augmented source domains and target domain. Since the augmentation $g(\cdot)$ could potentially affect both the two terms by changing the distribution of source domains, we here investigate how to design $g(\cdot)$ for reducing the two terms. To simplify the question, following [51, 49], we consider a feature extractor h with a linear softmax classifier $\textbf{W}=[\textbf{w}_{1},\textbf{w}_{2},...,\textbf{w}_{L}]\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels, and $L$ is the number of categories. For augmentation $g\sim\mathcal{G}$, the training risk on the augmented data is:
@@
Proof. To investigate how to design $g(\cdot)$ for reducing both $\varepsilon(D_{s},D_{s}^{aug})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$, following [51, 49], we explore the simplified problem in conjunction with a linear softmax classifier $\textbf{W}=[\textbf{w}_{1},\textbf{w}_{2},...,\textbf{w}_{L}]\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels, and $L$ is the number of categories. Let $\textbf{h}(z)=[h_{1}(z),h_{2}(z),\dots,h_{C}(z)]^{\top}$ represent the feature vector extracted by the model. Given the augmentation function $g\in\mathcal{G}:\mathcal{Z}\rightarrow\mathcal{Z}$, we could reformulate the risk $R(h_{aug};D_{s}^{aug})$ in Eq. (13) as:
@@
To narrow the generalization error bound in Eq. (14), we aim to reduce the domain gap $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ while keeping $\varepsilon(D_{s},D_{s}^{aug})$ small. According to Eq. (9), since the predictied label of the model can be denoted as $\textbf{W}^{\top}\textbf{h}$, we reformulate the domain discrepancy $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ as:
@@
Denoting the features being invariant across source and target domains as $h_{di}$, and the features being specific for source domains as $\textbf{h}_{ds}$. The above equation indicates that to reduce the domain gap, we should constrain the model to enhance the learning of the weight $w_{di}$ corresponding to $h_{di}$ while reducing the learning of $w_{ds}$ corresponding to $h_{ds}$. Inspired by previous works [23, 41, 40] that emphasize shape features as domain-invariant features, we assume that the extracted features h could be decomposed to global shape features $h_{gs}$, local edge features $h_{le}$ and style features $h_{st}$. The local edge features $h_{le}$ and style features $h_{st}$ mainly contain domain-specific information, while the global shape features $h_{gs}$ mainly encode domain-invariant information. To reduce $\varepsilon(D_{s},D_{s}^{aug})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$, the augmentation $g(\cdot)$ should adhere to:
@@
The above inequalities highlight the effectiveness of our SETA, which primarily perturbs local edge features $h_{le}$ and style features $h_{st}$ while preserving shape features $h_{gs}$. As a result, SETA could increase $\epsilon_{1}$ and $\epsilon_{2}$ while keeping $\xi$ small, thus effectively diminishing both the $\varepsilon(D_{s},D_{s}^{aug})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ during training. It is noteworthy that previous DG augmentation methods [13, 14, 19] mainly perturb style information (increasing $\epsilon_{2}$) and preserveing the other features (keeping $\xi$ and $\epsilon_{1}$ small), which reduces model sensitivity to style shifts. However, due to the lack of explicit constraints to learn shape information, these methods may inadvertently encourage the model to excessively learn domain-specific local edge information, which damages the generalization ability of the model. In contrast, our proposed method delves into local edge perturbations to increase $\epsilon_{1}$, potentially compelling the model to focus on shape representations. The two stylized variants of SETA could further increase the $\epsilon_{1}$ and $\epsilon_{2}$, thereby effectively reducing the generalization error bound of the model. The empirical evidences in Sec. 5 support the assertion that our method can significantly enhance the model learning of shape information while reducing the domain gap between source and target domains.
@@ IV-B Implementation Details
TABLE I: Performance (%) comparisons with state-of-the-art DG approaches on the DomainBed benchmark. We conduct a comprehensive evaluation by comparing our methods with $15$ existing DG algorithms across five diverse multi-domain datasets: PACS [54], VLCS [56], OfficeHome [55], TerraInc [57], and DomainNet [58]. The CNN-based methods in our experiments employ the ResNet-50 architecture, and we select two SOTA transformer-based models with similar parameter amounts to ResNet-$50$, i.e., Swin-T [59] and GFNet-H-Ti [18] for our methods. We use the validation set from source domains for model selection. The best is bold while the second-best is underlined.
@@ IV-C Compared with SOTA Methods
TABLE II: Ablation study (%) on different modules of SETA on the DomainBed benchmark. We denote the Energy-based Edge Tokens Selection as ETS and the Shape-disrupted Token Shuffling as STS, respectively. The experiments are conducted on GFNet-H-Ti.
Methods ETS STS PACS VLCS OfficeHome TerraInc DomainNet Avg.
IV-D Generalization on Semantic Segmentation.
Datasets and implementation details. To evaluate the cross-domain generalization ability of segmentation models, we adopt the synthetic dataset GTA5555 [70] for model training, while using the real-world datasets for testing, including CityScapes [71], BDD-$100$K [72], and Mapillary [73]. We follow the previous work [74] and adopt the DeepLabV3333+ [75] with ResNet50505050 as the backbone for the semantic segmentation architecture. The SGD optimizer with an initial learning rate of $21e-$ 2 and momentum of $0.9$ is utilized. For the quantitative experiments, the mean Intersection over Union (mIoU) is used to measure the segmentation performance of the model.
TABLE III: Comparisons (%) with the SOTA augmentation methods. The experiments are conducted on DeepLab-v3333 [75] segmentation network with ResNet50505050 as backbone. We train the model on synthetic GTA5555 while using real CityScapes, BDD-$100$K, and Mapillary for testing. The best is bold.
@@
Experimental results. We present the results in Tab. III and compare our SETA with SOTA augmentation-based DG methods. Specifically, we observe that existing augmentation methods can effectively enhance performance by perturbing style statistics, thereby improving the model’s robustness to style variations. However, since these methods ignore potentially spurious local edges in features, the model could still learn domain-specific information and overfit to the source domain. Differently, our SETA could effectively suppress spurious edge features, thus enhancing the model learning of semantic structure information. As shown in Tab. III, our SETA achieves state-of-the-art overall performance, yielding the baseline by $4.95\%$ ($32.37\%$ vs. $27.42\%$). The visualization results in Fig. 3 further prove the effectiveness of our method.

### v2

@@ III-C Enhancing Shape Bias by Texture Noise
While edge representations encompass both global shape edges and local texture edges, directly guiding the model to learn these representations could result in model’s sensitivity to local textures. To address this, we propose Shape-disrupted Token Shuffling (STS) module, which generates neural local edge noise by shuffling token positions from another sample, thereby disrupting category-related shape semantics. By randomly interpolating between the target edge features extracted by ETS and the shape-disrupted texture noise generated by STS, we could effectively preserve category-relevant shape information while perturbing category-irrelevant local textures.
Shape-disrupted token shuffling. Motivated by the insight that holistic shapes could be discerned through the relative positions of edge tokens [17, 21, 43], we introduce texture noise using shape-disrupted samples by shuffling tokens within the samples. For the input representation $z_{j}\in\mathbb{R}^{N\times C}$, we employ ${\rm Shuffle}(\cdot)$ function in the token dimension:
@@
where $\lambda$ is a random weight sampled from the Beta distribution ${\rm Beta}(\alpha,\beta)$ with $\alpha,\beta\in(0,\infty)$ being two hyper-parameters. The Mixup version of SETA can introduce bote style noise and local edge noise into the edge-related tokens, thereby encouraging the model to learn global semantic features from the edge representations. 2) For the CutMix version, we generate the token-level binary mask $\mathcal{M}_{\text{edge}}^{\lambda}\in\mathbb{R}^{N\times C}$ with the element being set to $1$ if the corresponding element in $A(\hat{z}_{i})$ is in the top $\lambda\sim{\rm Beta}(\alpha,\beta)$ percentage elements. The augmented representation $\widetilde{z}_{i}$ is:
$\widetilde{z}_{i}=\mathcal{M}_{\text{edge}}^{\lambda}\odot z_{i}+(1-\mathcal{M}_{\text{edge}}^{\lambda})\odot z_{j}^{s}.$ (8)
@@
Figure 3: An intuitive illustration of our SETA. The ETS module extracts edge-related tokens from the original sample, while the STS module generates shape-disrupted noise by randomly selecting another sample from the current batch. The Mixup version of SETA blends the values of the edge-related tokens and shape-disrupted tokens, while the CutMix version of SETA replaces the edge-irrelated tokens by the shape-disrupted tokens.
@@
In practice, for each mini-batch, we randomly select half of the samples to conduct the SETA augmentation. Within this process, the Mixup variant of SETA is applied with a probability of $p_{\text{mode}}$, while in other cases, the CutMix version of SETA is employed. An intuitive illustration of SETA is presented in Fig. 3. By introducing neural noise from random samples into original samples, domain-specific local edges are perturbed, while category-related shape information is preserved. This allows SETA to encourage the model to learn domain-invariant shape representations, thereby enhancing model generalization. During inference, SETA is closed as conventional augmentation-based DG methods [14, 19].
@@
Two stylized variants of SETA. To adequately enhance model robustness to domain shifts, we further explore two stylized versions of SETA. As our SETA primarily relies on texture noise extracted from shape-disrupted neural samples, the diversity of styles within the augmented samples is still limited. To address this limitation, we have incorporated two SOTA DG style augmentation methods and devised two variants of SETA, specifically SETA-S-DSU (combined with DSU [14]), which introduces statistical diversity, and SETA-S-ALOFT (combined with ALOFT [13]), which introduces frequency diversity. In this context, when generating token-shuffled samples with disrupted shapes, we also apply channel-level style augmentations (i.e., DSU or ALOFT) to these samples. The stylized variants could concurrently amplify the diversity of both style and texture edges, thus providing a more comprehensive simulation of potential domain shifts.
@@
In summary, we propose a token-level feature augmentation designed to encourage the shape bias of the model by introducing local edge noise. Note that our SETA diverges significantly from prior augmentation-based DG works [13, 14, 19, 31], which solely perturb style information while keeping other features unchanged. Since these existing methods ignore spurious local edges within the remaining features, the model inevitably learns domain-specific features and fails to capture shape information effectively. In contrast, our method proposes to interpolate edge-pass shape-preserved representations from the original image with shape-disrupted local edge representations from another image. The method incentivizes not only edge detection but also shape sensitivity: 1) classifying the edge tokens requires the model to exploit category-related edge information; 2) distinguishing the category of edge tokens from the mixed shuffled sample requires the model to distinguish the overall object shape from the shuffled sample edges. As a result, our method could enhance the model learning of class-relevant shape edges while suppressing harmful category-irrelevant local edges and style information. Besides, the computational overhead of SETA is minimal, only involving a few matrix operations. The components within SETA introduce no additional parameters or inference time, resulting in negligible training overhead, e.g., increasing only $0.6\%$ ($0.154s$ vs. $0.153s$) in time for each training step from the baseline on the GFNet-H-Ti architecture.
@@ III-D Theoretical Insight of SETA
Proof. Given the source dataset $D_{s}$, feature augmentation creates diverse data to augment the training data, denoted as $D_{s}^{aug}$. Recall that Lemma $1$ provides a generalization error bound between two different distributions. For $h_{aug}$ trained on the augmented source dataset $D_{s}^{aug}$, the following inequality holds with a probability of at least $1-\delta$:
@@
Theorem $1$ indicates that the upper bound of generalization risk on target domain depends primarily on three terms: 1) $\varepsilon(D_{s},D_{s}^{aug})$ that denotes the consistency between the risk of the model on clean and augmented data; 2) $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ that presents the distance between augmented source domains and target domain; 3) $\lambda$ that is the combined error of the optimal hypothesis $h^{*}$ on all domains. These terms are all influenced by two factors: 1) the model’s ability to extract invariant information between source and target domains; and 2) the augmentation to reduce the discrepancy between these domains. Regarding the first factor, we analyze it from the architectural perspective. For CNN models, which heavily rely on convolutions with limited receptive field, the feature response of convolution can be expressed as $\Phi_{\text{CNN}}(x)=W_{\text{conv}}*z$, where $W_{\text{conv}}$ is convolutional kernel, and * is convolution operation. Due to the high sensitivity of convolution to local textures, significant differences in $z$ between domains (e.g., varying textures) can lead to an increased discrepancy in $\Phi_{\text{CNN}}(z)$ between source and target domains. This increase contributes to a larger $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ and a greater $\lambda$. Therefore, CNNs perform inferiorly in domains with substantial gaps. In contrast to CNNs, token-based models (ViTs and MLPs) learn global dependencies among tokens, focusing more on global shape information rather than local textures. As a result, they are better at capturing shape information from different domains, thereby reducing domain gap $d_{H\Delta H}(D_{s},D_{t})$ and achieving a smaller $\lambda$ of the optimal hypothesis $h^{*}$. The above analysis indicates that the high dependency of CNNs on local convolution makes them more susceptible to domain shifts, whereas ViTs and MLPs, with their learning of global token dependencies, exhibit superior generalization ability. Therefore, using advanced architectures can effectively reduce the model’s generalization error bound.
Besides, we analyze the effects of feature augmentation on generalization error bound. Since the augmentation $g(\cdot)$ could potentially affect both the two terms by changing the distribution of source domains, we here investigate how to design $g(\cdot)$ for reducing the two terms. To simplify the question, following [46, 48], we consider a feature extractor h with a linear softmax classifier $\textbf{W}=[\textbf{w}_{1},\textbf{w}_{2},...,\textbf{w}_{L}]\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels, and $L$ is the number of categories. For augmentation $g\sim\mathcal{G}$, training risk on augmented data is:
@@
Proof. To investigate how to design $g(\cdot)$ for reducing both $\varepsilon(D_{s},D_{s}^{aug})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$, following [46, 48], we explore the simplified problem in conjunction with a linear softmax classifier $\textbf{W}=[\textbf{w}_{1},\textbf{w}_{2},...,\textbf{w}_{L}]\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels, and $L$ is the number of categories. Let $\textbf{h}(z)=[h_{1}(z),h_{2}(z),\dots,h_{C}(z)]^{\top}$ denote feature vector extracted by the model. Given the augmentation function $g\in\mathcal{G}:\mathcal{Z}\rightarrow\mathcal{Z}$, we reformulate the risk $R(h_{aug};D_{s}^{aug})$ in Eq. (13) as:
@@
To narrow the generalization error bound in Eq. (14), we aim to reduce the domain gap $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ while keeping $\varepsilon(D_{s},D_{s}^{aug})$ small. According to Eq. (9), since the predicted label of the model can be denoted as $\textbf{W}^{\top}\textbf{h}$, we reformulate the domain discrepancy $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ as:
@@
Denoting the features being invariant across source and target domains as $h_{di}$, and the features being specific for source domains as $h_{ds}$. The above equation indicates that to reduce the domain gap, we should constrain the model to enhance the learning of the weight $w_{di}$ corresponding to $h_{di}$ while reducing the learning of $w_{ds}$ corresponding to $h_{ds}$. Inspired by pioneering works [21, 22, 39] that divides features into style, texture, and shape, we formulate the hypothesis specifically for DG tasks, assuming that the extracted features h could be decomposed to global shape features $h_{gs}$, local edge features $h_{le}$ and style features $h_{st}$. The local edge features $h_{le}$ and style features $h_{st}$ mainly contain domain-specific information, while the global shape features $h_{gs}$ mainly encode domain-invariant information. To reduce $\varepsilon(D_{s},D_{s}^{aug})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$, the augmentation $g(\cdot)$ should adhere to:
@@
The above inequalities highlight the effectiveness of our SETA, which primarily perturbs local edge features $h_{le}$ and style features $h_{st}$ while preserving shape features $h_{gs}$. As a result, SETA could increase $\epsilon_{1}$ and $\epsilon_{2}$ while keeping $\xi$ small, thus effectively diminishing both the $\varepsilon(D_{s},D_{s}^{aug})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ during training. Theorem 1 and Proposition 1 jointly analyze the model’s generalization error bound from both the architectural and augmentation perspectives. Existing DG methods, which are predominantly based on CNNs, suffer the texture bias due to the limited receptive field of local convolutions, resulting in larger $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$ and $\lambda$. Besides, previous DG augmentation methods [13, 14, 19] are primarily designed for CNNs, focusing on perturbing channel-level style information while ignoring to suppress spatial-level local spurious edges. Due to the lack of explicit constraints to learn shape information, these methods may inadvertently encourage the model to excessively learn domain-specific local edge information, leading to larger $\varepsilon(D_{s},D_{s}^{aug})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(D_{s}^{aug},D_{t})$. These issues increase generalization error bound of existing CNN-based methods. In contrast, our proposed method delves into local edge perturbations to increase $\epsilon_{1}$, potentially compelling the model to focus on shape representations. The two stylized variants of SETA could further increase the $\epsilon_{1}$ and $\epsilon_{2}$, thereby effectively reducing the generalization error bound. The empirical evidence in Sec. 7 supports the assertion that our method can significantly enhance the model learning of shape information while reducing the domain gap between source and target domains.
@@ IV-B Implementation Details
TABLE I: Performance (%) comparisons with SOTA DG methods on DomainBed benchmark. We conduct a comprehensive evaluation by comparing our methods with $15$ existing DG algorithms across five diverse multi-domain datasets: PACS [51], VLCS [53], OfficeHome [52], TerraInc [54], and DomainNet [55]. The CNN-based methods in our experiments employ the ResNet-50 architecture, and we select two SOTA transformer-based models with similar parameter amounts to ResNet-$50$, i.e., Swin-T [56] and GFNet-H-Ti [18] for our methods. We use the validation set from source domains for model selection. The best is bold while the second-best is underlined.
@@ IV-C Compared with SOTA Methods
TABLE II: Comparisons (%) with the SOTA augmentation methods. The experiments are conducted on DeepLab-v3333 [68] segmentation network with ResNet50505050 as the backbone. We train the model on synthetic GTA5555 while using real CityScapes, BDD-$100$K, and Mapillary for testing. The best is bold.
@@
IV-D Generalization on Semantic Segmentation
Datasets and implementation details. To evaluate the cross-domain generalization ability of segmentation models, we adopt the synthetic dataset GTA5555 [69] for model training while using the real-world datasets for testing, including CityScapes [70], BDD-$100$K [71], and Mapillary [72]. We follow the previous work [73] and adopt the DeepLabV3333+ [68] with ResNet50505050 as the backbone for the semantic segmentation architecture. The SGD optimizer with an initial learning rate of $21e-$ 2 and momentum of $0.9$ is utilized. For the quantitative experiments, the mean Intersection over Union (mIoU) is used to measure the segmentation performance of the model.
Experimental results. We present the results in Tab. II and compare our SETA with SOTA augmentation-based DG methods. Specifically, we observe that existing augmentation methods can effectively enhance performance by perturbing style statistics, thereby improving the model’s robustness to style variations. However, since these methods ignore potentially spurious local edges in features, the model could still learn domain-specific information and overfit to the source domain. Differently, our SETA could effectively suppress spurious edge features, thus enhancing the model learning of semantic structure information. As shown in Tab. II, our SETA achieves state-of-the-art overall performance, yielding the baseline by $4.95\%$ ($32.37\%$ vs. $27.42\%$). The visualization results in Fig. 4 further prove the effectiveness of our method.